The mean weight of Grade 11 Section Star is 51 kg. and the standard deviation is 3.12 kg. Assuming the weight are normally distributed, what percent of weight greater than 57 kg.?

1
Expert's answer
2022-05-03T11:56:05-0400
"P(X>57)=1-P(X\le\dfrac{57-51}{3.1})=0.0265"

"2.65\%"